OpenStudy (nannie):

HELP! Use rational exponents to simplify. ⁴√x^20

OpenStudy (amistre64):

a radical index is an exponential denominator

OpenStudy (anonymous):

(x^20)^1/4 sso =x^5

OpenStudy (amistre64):

n.rt(A^m) = A^(m/n)

OpenStudy (nannie):

Answer- x^20/4 ???

OpenStudy (nannie):

or x^5??

OpenStudy (anonymous):

it's the same thing.